Mary is twice as old as Hilda. Four years from now the sum of their ages will be 41. Find their present age.
(side-note, I'm in grade 9 so my teacher expects me to use only ONE variable)
Thank you!

Found 2 solutions by ankor@dixie-net.com, mananth:
Answer by ankor@dixie-net.com(22740)

Mary is twice as old as Hilda.
let h = Hilda's age
then, since Mary is twice as old:
2h = Mary's age
Four years from now the sum of their ages will be 41.
(h+4) + (2h+4) = 41
h + 2h + 4 + 4 = 41
3h + 8 = 41
3h = 41 - 8
3h = 33
h = 33/3
h = 11 yrs old is Hilda's age
then
2*11 + 22 is Mary's age
:
:
Check their age total in 4 yrs
11 + 4 = 15
22 + 4 = 26
-----------
tot age 41

Answer by mananth(16946)

Mary is twice as old as Hilda. Four years from now the sum of their ages will be 41. Find their present age.

Let Hilda's age be x
Mary's age will be 2x
2x+4 +x+4 = 41
3x+8 =41
3x=41-8
3x =33
x=11
Hilda's age 11 years
Mary's age = 2x = 22 years